Am...Sir, can you please help me solve this age problem:
Half of Henry's age added to 1/3 of Daisy's age is 11 years.
Six years from now, the sum of their ages will be 40 years.
How old is each?

Answer by scott8148(6628)

"Half of Henry's age added to 1/3 of Daisy's age is 11 years" __ H/2+D/3=11 __ multiplying by 6 (LCD) __ 3H+2D=66

"Six years from now, the sum of their ages will be 40 years" __ (H+6)+(D+6)=40 __ subtracting 12 __ H+D=28 __ H=28-D

substituting __ 3(28-D)+2D=66 __ distributing __ 84-3D+2D=66 __ subtracting 84 __ -D=-18 __ dividing by -1 __ D=18

substituing __ H+18=28 __ subtracting 18 __ H=10